### v1

@@ V Neutron and proton EMT distributions
with the understanding that these expressions are to be evaluated with respectively $\rho_{p}(r)$, $\phi_{p}(r)$, $V_{p}^{0}(r)$, $A_{p}^{0}(r)$ in Eqs. (11a-f) in the proton case, and with $\rho_{n}(r)$, $\phi_{n}(r)$, $V_{0}^{n}(r)$ in Eqs. (15a-e) and the Coulomb potential set to zero in the neutron case. In Figs. 2a-d we show the EMT distributions as functions of $r$ for neutron in comparison to the proton results Varma:2020crx . In the region $r\lesssim 2\,\rm fm$, the proton and neutron distributions look much alike up to about $1\,\rm fm$. The neutron exhibits a slightly larger energy density and slightly stronger forces, appearing more compact compared to the slightly more spread out proton. These effects are small because in the inner region where the dust is located, the Coulomb force is overwhelmed by the strong forces.
@@
which reinforces what we saw in Fig. 2: in the neutron the forces are slightly larger than in the proton. In the neutron, the normal forces (7) are always positive. In the proton, they are positive up $1.88\,\rm fm$, see Fig. 2d and Table 1. As there is no dust at that distance, the latter this does not impact the mechanical stability of the proton Varma:2020crx .
@@
In the model, the reason for the divergence of the proton $D(t)$ in Eq. (19) is due to the large-$r$ behavior of the Coulomb field. The neutron EMT distributions decay exponentially at large $r$, see Table 1, and the integrals defining the $D$-term in Eq. (8) converge such that one obtains a well-defined and unambiguous result for $D_{\rm neut}$. In the proton case, however, the Coulomb field determines the asymptotics, see Table 1, and the corresponding integrals diverge. In Ref. Varma:2020crx it is was proposed to introduce a “regularized” proton $D$-term $D_{\rm prot,reg}$. In the remainder of this section, we will review how this regularization works and, based on the neutron results, show that it is reasonable. In Sec. VIII we will provide an argument why a regularized proton $D$-term makes sense also from a practical point of view.
@@ VII EMT radii and neutron size
The fact that the neutron appears systematically smaller than the proton accross the various radii (24-26) can be intuitively understood: the repulsive electrostatic forces present in the proton (but absent in the neutron) tend to spread out the matter distribution in the proton. The effect is relatively small because in the model the electromagnetic forces are three orders of magnitude weaker than strong forces. We recall that the strong forces in the model are the relatively weak residual forces. In nature, the QCD color forces are far stronger such that experimentally one should expect even smaller differences between the EMT radii of proton and neutron.
@@ VIII Simulating the physical situation for proton and neutron
It is an interesting question whether the divergence for $t\to 0$ of the $D(t)$ of proton can be observed experimentally. The model can be used to shed light on this question but for that one needs first to make the model description more realistic. For that we will address two aspects. (i) In the model the proton charge radius is about $15\,\%$ smaller than the experimental value. (ii) The neutron and regularized proton $D$-terms are about an order of magnitude smaller than, e.g., the $D$-terms in the more realistic chiral quark soliton model Goeke:2007fp or in lattice QCD Hackett:2023rif .
@@
The neutron stress tensor properties scale homogeneously with $\lambda_{m}$ and $\lambda_{g}$ (in a specific way depending on the property). In the the proton case, the stress tensor properties scale homogeneously with $\lambda_{m}$ (because all the relevant dimensionful parameters are rescaled including the photon mass, $m_{\gamma}^{\prime}=\lambda_{m}\,m_{\gamma}$, which happens to be zero) but not with $\lambda_{g}$ (because in Eq. (27) we do not rescale simultaneously the electromagnetic coupling constant and keep it fixed). Due to the smallness of the electromagnetic coupling constant there is, however, an approximate scaling with $\lambda_{g}$. [^2] The energy density has no definite scaling behavior because the model parameter $m_{\rm dust}$ is adjusted to reproduce the proton mass according to footnote 1 as we vary the $\lambda_{i}$. As a consequence of that, the energy density can turn negative in the vicinity of $r=0$ for $\lambda_{g}\geq 2.1$ when $\lambda_{m}=0.85$. This may be considered a “finite-size renormalization artifact” in an interacting theory as coupling constants are varied. In practice, renormalization effects do not always preserve all physical properties, like positivity of energy density in this case. As $m_{\rm dust}$ enters only in combination with $\rho(r)$ (normalized to unity) in $T_{00}(r)$, it is just an additive constant in the nucleon mass and does not affect mechanical stability and properties related to the stress tensor.
In this way, in the model the nucleon properties become functions of $\lambda_{m}$ and $\lambda_{g}$. As can be seen from the Yukawa tails of the meson fields proportional to $\exp(-m_{i}r)$, Eqs. (11c11d15b15c), one can think of the change $m_{i}\to m_{i}^{\prime}$ in Eq. (27) as implying a redefinition of the unit of length. As a consequence, for instance the mean square square root dust radius (which corresponds to the proton charge radius scales) depends on $\lambda_{m}$ as
@@
Fig. 4b shows the same result as Fig. 4a but with a logarithmic scale for the (t)(-t)–axis to illustrate when one may expect to start seeing deviations. Until $(-t)\approx 10^{-4}\,\rm GeV^{2}$ there is no appreciable difference between the two curves. In other words, until then the proton and neutron form factors $D(t)$ may be expected to look basically the same. A noticible difference between the $D(t)$ of these two baryons becomes clearly visible only below $(-t)\approx 10^{-5}\,\rm GeV^{2}$. At still smaller (t)(-t), somewhere in the range $10^{-6}\,\rm GeV^{2}<(-t)<10^{-5}\,\rm GeV^{2}$ depending on the parameter $\lambda_{g}$, the proton $D(t)$ changes sign and becomes positive.
@@
Thus, the model exactly reproduces the QED prediction in the limit $t\to 0$ and lattice QCD in the region for $0.06\,\rm GeV^{2}<t\lesssim 1\,\rm GeV^{2}$ which gives confidence that it can reasonably interpolate between what we know from QED and lattice QCD about the $D(t)$ form factor of the proton. Based on our model results, we can draw two conclusions. First, the proton and neutron $D(t)$ can be expected to be practically the same down to $(-t)\approx 10^{-4}\,\rm GeV^{2}$. Second, it does make sense to introduce the notion of a regularized proton $D$-term to approximate the exact proton $D(t)\approx D_{\rm prot,reg}/(1-t/m_{D}^{2})^{n}$ with some effective paramater $n$, for instance, $n=2$ as used in the lattice data fit in Ref. Hackett:2023rif or $n=3$ as used in Varma:2020crx or in the first attempt to extract the proton $D(t)$ from data Nature .
@@ IX Conclusions
The EMT form factor $D(t)$ of neutral hadrons is negative and finite at $t=0$, while that of charged hadrons changes sign at small values of (t)(-t) and exhibits for $t\to 0$ a divergence $D(t)\propto 1/\sqrt{-t}$ as first observed for charged pions in chiral pertubation theory Kubis:1999db and for proton Varma:2020crx in the model by Białynicki-Birula BialynickiBirula:1993ce . The divergence of $D(t)$ is a universal effect for any charged particle due to the masslessness of the photon Donoghue:2001qc ; Metz:2021lqv ; Freese:2022jlu . In this work, we investigated the important question whether this effect can be observed experimentally for the proton, and whether protons and neutrons must be treated differently in phenomenological studies of hard exclusive reactions.
@@
To examine this assertion quantitatively it was necessary to alter the original proton model formulation, where the proton charge radius is about $15\,\%$ too small (though “not completely out of touch with reality” BialynickiBirula:1993ce ) and the $D$-term is about an order of magnitude smaller Varma:2020crx compared to results from more realistic hadronic models and lattice QCD (because the residual nuclear forces underlying the mode are far weaker than the strong forces between quarks). To make the modelling of the strong forces more realistic, we introduced two model parameters: one to bring the model result for the proton charge radius in agreement with experiment, and the other to bring the model result for the $D$-term in agreement with with lattice data Hackett:2023rif . Remarkably, this resulted in a very good description of lattice QCD data on $D(t))$ up to $(-t)\lesssim{\rm 1\,GeV^{2}}$. We have also analytically proven that the model exactly reproduces the QED divergence of $D(t)$ for the proton.
@@
Equipped with this realistic model incorporating all that is currently known about $D(t)$ from QED and QCD from $t\to 0$ until $(-t)\lesssim{\rm 1\,GeV^{2}}$, we have shown that one would need to go to (t)(-t) below $10^{-4}\,\rm GeV^{2}$ to observe deviations between the proton and neutron $D(t)$ form factors. In order to verify experimentally the QED prediction for the proton $D(t)=\alpha\,\pi\,M/(4\sqrt{-t})+\dots$ one would need to reach (t)(-t) values of $10^{-8}\,\rm GeV^{2}$ or smaller. In the very first attempt to extract information related to $D(t)$ from Jefferson Lab data the minimal (t)(-t) value was $(-t)_{\rm min}=0.09\,\rm GeV^{2}$ Nature . Impact studies for the planned Electron-Ion Collider assume $(-t_{\rm min})=0.03\,{\rm GeV}^{2}$ Aschenauer:2013hhw ; AbdulKhalek:2021gbh . Our study therefore leads to the conclusion that in foreseeable future, the experimental observation of QED effects in the proton $D(t)$ will be out of reach. This means that, for all practical purposes in phenomenological studies of hard exclusive reactions, the proton and neutron $D(t)$ will look the same: negative and finite down to the lowest measurable (t)(-t) values in ongoing and planned experiments. This implies that it is a reasonable step to introduce the notion of a “regularized proton $D$-term” $D_{\rm prot,reg}$ for Ansätze of the type $D(t)\approx D_{\rm prot,reg}/(1-t/m_{D}^{2})^{n}$ for the proton for $(-t)>10^{-4}\,\rm GeV^{2}$ in phenomenological applications and which can be expected to be very similar to the neutron $D$-term. In fact, in the model $D_{\rm neut}$ and $D_{\rm prot,reg}$ coincide within 1 ‰ accuracy.
@@
We generalized the method to regularize the proton $D$-term of Ref. Varma:2020crx to other EMT properties, and explored it to compute the EMT radii of proton which are divergent (similarly to the $D$-term) due to the long-range Coulomb field contribution. Comparing the echanical radius and the mass radius of the proton to those of the neutron (which are finite), we observed that the neutron is systematically slightly smaller than the proton because latter appears slightly swollen compared to the neutron, due to Coulomb repulsion. As a by-product of this study we point out that the EMT radii allow, in contrast to the mean square charge radius, the neutron size. According to the model, proton and neutron will appear as nearly the same size — as one would intuitively expect for these SU(2) partners which are very similar in so many respects, if one refrains from electromagnetic corrections and isospin breaking effects. Based on measurements of the hard exclusive reactions, it will be possible to test this prediction experimentally.

### v2

@@ V Neutron and proton EMT distributions
with the understanding that these expressions are to be evaluated with respectively $\rho_{p}(r)$, $\phi_{p}(r)$, $V_{p}^{0}(r)$, $A_{p}^{0}(r)$ in Eqs. (11a-e) in the proton case, and with $\rho_{n}(r)$, $\phi_{n}(r)$, $V_{0}^{n}(r)$ in Eqs. (15a-d) and the Coulomb potential set to zero in the neutron case. In Figs. 2a-d we show the EMT distributions as functions of $r$ for neutron in comparison to the proton results Varma:2020crx . In the region $r\lesssim 2\,\rm fm$, the proton and neutron distributions look much alike up to about $1\,\rm fm$. The neutron exhibits a slightly larger energy density and slightly stronger forces, appearing more compact compared to the slightly more spread out proton. These effects are small because in the inner region where the dust is located, the Coulomb force is overwhelmed by the strong forces.
@@
which reinforces what we saw in Fig. 2: in the neutron the forces are slightly larger than in the proton. In the neutron, the normal forces (7) are always positive. In the proton, they are positive up $1.88\,\rm fm$, see Fig. 2d and Table 1. As there is no dust at that distance, this does not impact the mechanical stability of the proton Varma:2020crx .
@@
In the model, the reason for the divergence of the proton $D(t)$ in Eq. (19) is due to the large-$r$ behavior of the Coulomb field. The neutron EMT distributions decay exponentially at large $r$, see Table 1, and the integrals defining the $D$-term in Eq. (8) converge such that one obtains a well-defined and unambiguous result for $D_{\rm neut}$. In the proton case, however, the Coulomb field determines the asymptotics, see Table 1, and the corresponding integrals diverge. In Ref. Varma:2020crx it is was proposed to introduce a “regularized” proton $D$-term $D_{\rm prot,reg}$. In the remainder of this section, we will review how this regularization works and, based on the neutron results, show that it is reasonable. In Sec. VIII we will provide an argument as to why a regularized proton $D$-term makes sense also from a practical point of view.
@@ VII EMT radii and neutron size
The fact that the neutron appears systematically smaller than the proton across the various radii (24-26) can be intuitively understood: the repulsive electrostatic forces present in the proton (but absent in the neutron) tend to spread out the matter distribution in the proton. The effect is relatively small because in the model the electromagnetic forces are three orders of magnitude weaker than strong forces. We recall that the strong forces in the model are the relatively weak residual forces. In nature, the QCD color forces are far stronger such that experimentally one should expect even smaller differences between the EMT radii of proton and neutron.
@@ VIII Simulating the physical situation for proton and neutron
An interesting question is whether the divergence for $t\to 0$ of the $D(t)$ of proton can be observed experimentally. The model can be used to shed light on this question but for that one first needs to make the model description more realistic. For that we will address two aspects. (i) In the model the proton charge radius is about $15\,\%$ smaller than the experimental value. (ii) The neutron and regularized proton $D$-terms are about an order of magnitude smaller than, e.g., the $D$-terms in the more realistic chiral quark soliton model Goeke:2007fp or in lattice QCD Hackett:2023rif .
@@
The neutron stress tensor properties scale homogeneously with $\lambda_{m}$ and $\lambda_{g}$ (in a specific way depending on the property). In the proton case, the stress tensor properties scale homogeneously with $\lambda_{m}$ (because all the relevant dimensionful parameters are rescaled including the photon mass, $m_{\gamma}^{\prime}=\lambda_{m}\,m_{\gamma}$, which happens to be zero) but not with $\lambda_{g}$ (because in Eq. (27) we do not simultaneously rescale the electromagnetic coupling constant, keeping it fixed). Due to the smallness of the electromagnetic coupling constant there is, however, an approximate scaling with $\lambda_{g}$. [^2] The energy density has no definite scaling behavior because the model parameter $m_{\rm dust}$ is adjusted to reproduce the proton mass according to footnote 1 as we vary the $\lambda_{i}$. As a consequence of that, the energy density can turn negative in the vicinity of $r=0$ for $\lambda_{g}\geq 2.1$ when $\lambda_{m}=0.85$. This may be considered a “finite-size renormalization artifact” in an interacting theory as coupling constants are varied. In practice, renormalization effects do not always preserve all physical properties, like positivity of energy density in this case. As $m_{\rm dust}$ enters only in combination with $\rho(r)$ (normalized to unity) in $T_{00}(r)$, it is just an additive constant in the nucleon mass and does not affect mechanical stability and properties related to the stress tensor.
In this way, in the model the nucleon properties become functions of $\lambda_{m}$ and $\lambda_{g}$. As can be seen from the Yukawa tails of the meson fields proportional to $\exp(-m_{i}r)$, Eqs. (11c11d15b15c), one can think of the change $m_{i}\to m_{i}^{\prime}$ in Eq. (27) as implying a redefinition of the unit of length. As a consequence, for instance the mean square square root dust radius (which corresponds to the proton charge radius) depends on $\lambda_{m}$ as
@@
Fig. 4b shows the same result as Fig. 4a but with a logarithmic scale for the (t)(-t)–axis to illustrate when one may expect to start seeing deviations. Until $(-t)\approx 10^{-4}\,\rm GeV^{2}$ there is no appreciable difference between the two curves. In other words, until then the proton and neutron form factors $D(t)$ may be expected to look basically the same. A noticeable difference between the $D(t)$ of these two baryons becomes clearly visible only below $(-t)\approx 10^{-5}\,\rm GeV^{2}$. At still smaller (t)(-t), somewhere in the range $10^{-6}\,\rm GeV^{2}<(-t)<10^{-5}\,\rm GeV^{2}$ depending on the parameter $\lambda_{g}$, the proton $D(t)$ changes sign and becomes positive.
@@
Thus, the model exactly reproduces the QED prediction in the limit $t\to 0$ and lattice QCD in the region for $0.06\,\rm GeV^{2}<t\lesssim 1\,\rm GeV^{2}$ which gives confidence that it can reasonably interpolate between what we know from QED and lattice QCD about the $D(t)$ form factor of the proton. Based on our model results, we can draw two conclusions. First, the proton and neutron $D(t)$ can be expected to be practically the same down to $(-t)\approx 10^{-4}\,\rm GeV^{2}$. Second, it does make sense to introduce the notion of a regularized proton $D$-term to approximate the exact proton $D(t)\approx D_{\rm prot,reg}/(1-t/m_{D}^{2})^{n}$ with some effective parameter $n$, for instance, $n=2$ as used in the lattice data fit in Ref. Hackett:2023rif or $n=3$ as used in Varma:2020crx or in the first attempt to extract the proton $D(t)$ from data Nature .
@@ IX Conclusions
The EMT form factor $D(t)$ of neutral hadrons is negative and finite at $t=0$, while that of charged hadrons changes sign at small values of (t)(-t) and exhibits for $t\to 0$ a divergence $D(t)\propto 1/\sqrt{-t}$ as first observed for charged pions in chiral perturbation theory Kubis:1999db and for proton Varma:2020crx in the model by Białynicki-Birula BialynickiBirula:1993ce . The divergence of $D(t)$ is a universal effect for any charged particle due to the masslessness of the photon Donoghue:2001qc ; Metz:2021lqv ; Freese:2022jlu . In this work, we investigated the important question whether this effect can be observed experimentally for the proton, and whether protons and neutrons must be treated differently in phenomenological studies of hard exclusive reactions.
@@
To examine this assertion quantitatively it was necessary to alter the original proton model formulation, where the proton charge radius is about $15\,\%$ too small (though “not completely out of touch with reality” BialynickiBirula:1993ce ) and the $D$-term is about an order of magnitude smaller Varma:2020crx compared to results from more realistic hadronic models and lattice QCD (because the residual nuclear forces underlying the model are far weaker than the strong forces between quarks). To make the modelling of the strong forces more realistic, we introduced two model parameters: one to bring the model result for the proton charge radius in agreement with experiment, and the other to bring the model result for the $D$-term in agreement with lattice data Hackett:2023rif . Remarkably, this resulted in a very good description of lattice QCD data on $D(t)$ up to $(-t)\lesssim{\rm 1\,GeV^{2}}$. We have also analytically proven that the model exactly reproduces the QED divergence of $D(t)$ for the proton.
@@
Equipped with this realistic model incorporating all that is currently known about $D(t)$ from QED and QCD from $t\to 0$ until $(-t)\lesssim{\rm 1\,GeV^{2}}$, we have shown that one would need to go to (t)(-t) below $10^{-4}\,\rm GeV^{2}$ to observe deviations between the proton and neutron $D(t)$ form factors. In order to verify experimentally the QED prediction for the proton $D(t)=\alpha\,\pi\,M/(4\sqrt{-t})+\dots$ one would need to reach (t)(-t) values of $10^{-8}\,\rm GeV^{2}$ or smaller. In the very first attempt to extract information related to $D(t)$ from Jefferson Lab data the minimal (t)(-t) value was $(-t)_{\rm min}=0.09\,\rm GeV^{2}$ Nature . Impact studies for the planned Electron-Ion Collider assume $(-t_{\rm min})=0.03\,{\rm GeV}^{2}$ Aschenauer:2013hhw ; AbdulKhalek:2021gbh . Our study therefore leads to the conclusion that, in the foreseeable future, the experimental observation of QED effects in the proton $D(t)$ will be out of reach. This means that, for all practical purposes in phenomenological studies of hard exclusive reactions, the proton and neutron $D(t)$ will look the same: negative and finite down to the lowest measurable (t)(-t) values in ongoing and planned experiments. This implies that it is a reasonable step to introduce the notion of a “regularized proton $D$-term” $D_{\rm prot,reg}$ for Ansätze of the type $D(t)\approx D_{\rm prot,reg}/(1-t/m_{D}^{2})^{n}$ for the proton for $(-t)>10^{-4}\,\rm GeV^{2}$ in phenomenological applications which can be expected to be very similar to the neutron $D$-term. In fact, in the model $D_{\rm neut}$ and $D_{\rm prot,reg}$ coincide within 1 ‰ accuracy.
@@
We generalized the method to regularize the proton $D$-term of Ref. Varma:2020crx to other EMT properties, and explored it to compute the EMT radii of proton which are divergent (similarly to the $D$-term) due to the long-range Coulomb field contribution. Comparing the mechanical radius and the mass radius of the proton to those of the neutron (which are finite), we observed that the neutron is systematically slightly smaller than the proton because latter appears slightly swollen compared to the neutron, due to Coulomb repulsion. As a by-product of this study, we point out that the EMT radii allow, in contrast to the mean square charge radius, to characterize the neutron size. According to the model, proton and neutron will appear as nearly the same size — as one would intuitively expect for these SU(2) partners which are very similar in so many respects, if one refrains from electromagnetic corrections and isospin breaking effects. Based on measurements of the hard exclusive reactions, it will be possible to test this prediction experimentally.